# A Unified Debiasing Approach for Vision-Language Models across Modalities and Tasks

### v1

@@ 4.1 Selective Feature Imputation for Debiasing
For the debiasing representation $Z_{D}$, Selective Feature Imputation for Debiasing (SFID) uses a RandomForest Breiman (2001) $f$ to predict the sensitive attribute $y_{D}$. Given the interpretability of RandomForest, it is capable of providing feature importance for predictions, allowing us to identify which features in the frozen representation are relevant to the sensitive attribute. Additionally, RandomForest is known for not requiring hyperparameter tuning Probst et al. (2019) and for its computational efficiency Biau (2012), making it an easily implementable choice. As the objective of SFID is to lead VLMs’ components to produce a fair outcome, free of bias regarding a sensitive attribute, SFID prunes the important features that show higher relevance to the sensitive attribute. This procedure is beneficial as it considers the dependency between features, whereas methods like CLIP-clip Wang et al. (2021) extract feature importance by measuring the mutual information between each feature and the sensitive attribute, assuming each feature is independent.
However, simply dropping features cannot maintain the dimensionality of the representation, which is crucial for using the embedding for generation tasks. For example, in CLIP-CAP Mokady et al. (2021), the decoder (GPT-2 Radford et al. (2019)) takes the image representation from the CLIP ViT-B/32 Radford et al. (2021) image encoder as input, and the input dimension of GPT-2 is fixed. Therefore, we must maintain the dimension of the representation after pruning to utilize the pre-trained decoder but approaches like filling with zero-values or Gaussian noise may mislead the semantic meaning of the representation, as described in Figure 4 and ablation study in Section 5.3.
@@
We visualize how the important features are correlated to social biases by showing GradCAM Selvaraju et al. (2017) visualizations, as presented in Figure 2. For example, the more important features highlight human faces, while the least important features are correlated with the image background. SFID imputes the representations at important indices with the values from low-confidence samples, making the face-related features ambiguous. Despite this imputation, the replaced values remain within the distribution of the original samples.
@@
4.2 Comparison with Other Debiasing Approaches
As described in Section 1, two existing methods, DeAR Seth et al. (2023) and CLIP-clip Wang et al. (2021), utilize debiasing datasets with gender labels and pre-trained models to debias embeddings. While DeAR is specifically designed for image encoders and CLIP-clip for both image and text encoders, their methodologies could potentially be extended to text and image decoders, as well as other encoders. In contrast, Prompt-Debias Chuang et al. (2023) is designed primarily for text encoders, using pre-defined text prompts to achieve debiasing. These three approaches are selected for comparison in our analysis.
4.2.1 Debiasing with Additive Residuals (DeAR)
DeAR Seth et al. (2023) involves a two-step training process. First, it trains an attribute classifier, $h_{c}$ using the frozen representation by minimizing the cross-entropy loss. Second, it trains an linear transformation $h_{a}$ for $Z_{D}$ to produce an additive feature to make a debiased representation $Z_{a}$, i.e., $Z_{a}=h_{a}(Z_{D})+Z_{D}$. The objective function for optimizing $h_{a}$
indicating the need to maintain the representation’s meaning while guiding the linear transformation $h_{a}$ to minimize the maximum probability of $h_{c}$ and maximize the cross-entropy loss of $h_{c}$. The goal is for the additive feature to make the representation ambiguous, placing it near the attribute classifier’s decision boundary. However, DeAR does not consider feature importance and applies adversarial training to all feature columns. Moreover, it requires several hyperparameters for the adversarial training, such as learning rate, weight decay, and the weights between loss functions ($\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$). This complexity may lead to inefficiencies in adversarial training, which is sensitive to hyperparameter settings, resulting in less effective translation of the feature, as shown in Figure 5. For DeAR training, all hyperparameters are set according to Seth et al. (2023).
4.2.2 CLIP-clip
CLIP-clip Wang et al. (2021) aims to identify features related to the sensitive attribute by measuring the mutual information between each column feature and the sensitive attribute. However, this approach assumes that each column independently contains meaningful information about the sensitive attribute, which can be overly restrictive. The dependency between feature maps in the neural networks’ representation output is well-documented Bengio et al. (2013); Sui et al. (2021), supporting SFID’s capability in more effective feature selection.
Moreover, CLIP-clip is designed to clip the representation, reducing the dimension size of the embedding. The reduced embedding is not applicable as input for the decoder, which necessitates imputing the pruned features. Since CLIP-clip does not provide any suggestions for the imputed values, we adopted zero-value imputation across our experiments for CLIP-clip, where the number of pruned feature is set as $k=60$, showing best accuracy-fairness trade-off described in Appendix C.
4.2.3 Prompt-Debias
Prompt-Debias Chuang et al. (2023) seeks to mitigate bias in text embeddings by projecting out biased directions from the text input. However, Prompt-Debias is specifically designed for text-based tasks such as zero-shot classification and text-to-image generation. Unlike SFID, it lacks the flexibility to be applied across other components, limiting its effectiveness. This results in a performance constraint, as bias amplification can occur in the decoder or image encoder, components that Prompt-Debias cannot address. We include Prompt-Debias in our experiments for zero-shot classification, text-to-image retrieval, and text-to-image generation, excluding image captioning.
4.3 High-Confidence Strategy in Text-to-Image Generation
@@
In SFID, we utilize low-confidence samples to obtain imputation values that mute bias-related information within the embedding. However, in text-to-image generation, users may sometimes wish to specify the gender of the generated image through gender-specific prompts, such as “a photo of a man who works as a nurse." In these cases, rather than muting the bias-related features, we can impute features from samples classified with high confidence by RandomForest to retain the desired gender. For text-to-image generation, we report results using both low-confidence (LC) imputation and high-confidence (HC) imputation approaches, respectively.
@@
Model FACET Zero-shot Multi-class Classification
@@
Model Flickr30K
@@
Male Female Overall Composite
CLIP-CAP Baseline 34.57$\pm$0.83 25.41$\pm$0.73 1.53$\pm$0.80 3.54$\pm$1.94 2.10$\pm$0.70 3.24$\pm$1.61
DeAR 33.90$\pm$0.91 24.73$\pm$0.63 2.55$\pm$1.05 3.58$\pm$2.19 2.93$\pm$0.98 3.53$\pm$1.30
CLIP-clip 32.28$\pm$0.72 23.44$\pm$0.65 0.94$\pm$0.65 4.60$\pm$2.39 2.00$\pm$0.90 4.34$\pm$2.48
SFID 32.08$\pm$0.78 23.74$\pm$0.69 1.58$\pm$0.94 3.39$\pm$2.34 2.07$\pm$1.03 3.12$\pm$1.82
BLIP Baseline 24.01$\pm$0.62 17.06$\pm$0.60 0.73$\pm$0.59 2.32$\pm$1.46 1.15$\pm$0.65 2.26$\pm$1.26
DeAR 21.76$\pm$0.59 15.51$\pm$0.47 0.29$\pm$0.29 2.81$\pm$2.01 1.07$\pm$0.63 2.84$\pm$2.13
CLIP-clip 23.74$\pm$0.54 16.96$\pm$0.54 0.54$\pm$0.40 2.73$\pm$1.81 1.15$\pm$0.65 2.59$\pm$1.81
SFID 23.38$\pm$0.49 16.74$\pm$0.55 0.78$\pm$0.64 1.93$\pm$1.47 0.92$\pm$0.53 1.88$\pm$1.31
@@ 5.1 Implementation Detail
We use the FairFace dataset Kärkkäinen and Joo (2019) for image inputs and the Bias-in-Bios dataset De-Arteaga et al. (2019) for text inputs as our debiasing datasets, denoted as $X_{D}$. Each dataset is split into training and validation sets. A RandomForest classifier is applied to a 2D-shaped embedding representing all training samples, with low-confidence samples selected from the validation set. All hyperparameters and model settings for each baseline follow the default configurations provided in their respective open-source repositories. Detailed experimental settings, along with evaluation metrics and query datasets, are described in Section 3.
For a fair comparison, we conduct 10 experiments using different subsets and report the mean and standard deviation for the text-to-image retrieval task. For zero-shot classification and image captioning, we employ 1000 bootstrapping iterations to calculate the confidence intervals. For text-to-image generation, images are generated using 10 random seeds. The mismatch rates are reported as the mean and standard deviation over 10 runs, while the Skew for the Neutral prompt is reported as a single value based on 10 runs. Further analysis of this evaluation metric can be found in Appendix D.
@@ 5.2 Result Analysis
Model Mismatch Rate (Gender prompt) Neutral prompt
Male Female Overall Composite $Skew$
SDXL Baseline 0.48$\pm$0.62 4.22$\pm$2.36 2.35$\pm$1.22 4.42$\pm$2.57 83.25
DeAR 0.00$\pm$0.00 89.28$\pm$2.08 44.64$\pm$1.04 99.81$\pm$2.33 99.88
CLIP-clip 0.24$\pm$0.51 3.98$\pm$1.89 2.11$\pm$1.03 4.31$\pm$2.06 82.05
Prompt-Debias 62.41$\pm$3.58 22.65$\pm$5.79 42.53$\pm$3.85 58.49$\pm$3.64 82.77
SFID (LC) 0.12$\pm$0.38 1.81$\pm$0.63 0.96$\pm$0.42 1.97$\pm$0.67 81.57
SFID (HC) 1.33$\pm$1.55 0.36$\pm$0.58 0.84$\pm$0.71 1.74$\pm$1.57 81.57
CoDi Baseline 3.86$\pm$2.53 7.23$\pm$2.27 5.54$\pm$2.08 6.85$\pm$2.16 84.94
DeAR 2.65$\pm$1.59 8.19$\pm$2.53 5.42$\pm$1.10 8.05$\pm$3.00 86.14
CLIP-clip 2.65$\pm$1.59 7.35$\pm$1.65 5.00$\pm$1.39 7.01$\pm$1.53 84.58
Prompt-Debias 52.05$\pm$3.84 31.93$\pm$3.42 41.99$\pm$2.57 46.77$\pm$3.43 81.57
SFID (LC) 2.77$\pm$1.61 6.51$\pm$1.52 4.64$\pm$1.17 6.22$\pm$1.48 82.17
SFID (HC) 4.94$\pm$1.44 0.24$\pm$0.51 2.59$\pm$0.90 5.38$\pm$1.44 82.77
@@ Appendix B Evaluation Metric for Image Captioning
METEOR measures the balance between precision and recall of n-grams in generated captions, incorporating synonyms. Let $P$ and $R$ be the precision and recall of matches between the generated caption and ground truth, including exact, synonym, and paraphrase matches: $\text{METEOR}=F_{\text{mean}}\cdot(1-\text{Pen})$ where $F_{\text{mean}}=\frac{10\cdot P\cdot R}{R+9\cdot P}$ is a harmonic mean, and $\text{Pen}=0.5\times\left(\frac{\text{number of chunks}}{\text{number of matches}}\right)^{3}$ is a penalty term where a chunk is a set of contiguous words in the generated captions that are in the reference. SPICE Anderson et al. (2016) focuses on the semantic content of captions by comparing sets of propositional semantic tuples extracted from candidate and reference captions. SPICE is the F1 score of precision and recall between the tuples of generated captions and ground truth.
Appendix C The Impact of $k$ in CLIP-clip
@@
CLIP-clip Wang et al. (2021) requires a hyperparameter $k$ to determine the number of pruned features. We selected $k$ for CLIP-clip based on empirical results with XVLM, as shown in Figures 7 and 8. The value $k=60$ performs best in text-to-image retrieval for the Flickr30K dataset, compromising slightly on accuracy and recall. Meanwhile, SFID demonstrates the best trade-off at $k=50$, corresponding to Appendix A.1.
@@
Appendix D Confidence Interval in Text-to-Image Generation
@@
On the other hand, a metric used in Chuang et al. (2023) measures gender discrepancy as follows:
@@
Appendix E Multiple Sensitive Attribute
@@
To mitigate this bias, SFID demonstrates its effectiveness even though the attributes in the training set and test set do not exactly match, i.e. race in FairFace and skin tone in FACET. We adopted an evaluation metric inspired by Foulds et al. (2020) and Denis et al. (2021), which measures the maximum discrepancy across the sensitive attributes for each class. This metric is defined as follows:
@@
Appendix F Computational Resource

### v2

@@ 4.1 Selective Feature Imputation for Debiasing
For the debiasing representation $Z_{D}$, Selective Feature Imputation for Debiasing (SFID) uses a RandomForest [8] $f$ to predict the sensitive attribute $y_{D}$. Given the interpretability of RandomForest, it is capable of providing feature importance for predictions, allowing us to identify which features in the frozen representation are relevant to the sensitive attribute. Additionally, RandomForest is known for not requiring hyperparameter tuning [29] and for its computational efficiency [7], making it an easily implementable choice. As the objective of SFID is to lead VLMs’ components to produce a fair outcome, free of bias regarding a sensitive attribute, SFID prunes the important features that show higher relevance to the sensitive attribute. This procedure is beneficial as it considers the dependency between features, whereas methods like CLIP-clip [38] extract feature importance by measuring the mutual information between each feature and the sensitive attribute, assuming each feature is independent.
However, simply dropping features cannot maintain the dimensionality of the representation, which is crucial for using the embedding for generation tasks. For example, in CLIP-CAP [27], the decoder (GPT-2 [30]) takes the image representation from the CLIP ViT-B/32 [31] image encoder as input, and the input dimension of GPT-2 is fixed. Therefore, we must maintain the dimension of the representation after pruning to utilize the pre-trained decoder but approaches like filling with zero-values or Gaussian noise may mislead the semantic meaning of the representation, as described in Figure 4 and ablation study in Section 5.3.
@@
We visualize how the important features are correlated to social biases by showing GradCAM [33] visualizations, as presented in Figure 2. For example, the more important features highlight human faces, while the least important features are correlated with the image background. SFID imputes the representations at important indices with the values from low-confidence samples, making the face-related features ambiguous. Despite this imputation, the replaced values remain within the distribution of the original samples.
@@
4.2 High-Confidence Imputation in Text-to-Image Generation
@@
In SFID, we utilize low-confidence samples to obtain imputation values that mute bias-related information within the embedding. However, in text-to-image generation, users may sometimes wish to specify the gender of the generated image through gender-specific prompts, such as “a photo of a man who works as a nurse." In such cases, rather than suppressing bias-related features, we impute features from samples that are classified with high confidence as belonging to the specified gender, using a RandomForest classifier, to retain the desired gender attributes. For text-to-image generation, we report results using both low-confidence (LC) and high-confidence (HC) imputation approaches.
4.3 Dealing with Multiple Sensitive Attribute
We extend our approaches to address more complex bias scenarios in VLMs, focusing on multiple sensitive attributes. Specifically, we conduct additional experiments that examine racial bias by considering more than two sensitive attributes, enabling us to capture a broader spectrum of bias patterns. Since RandomForest can accommodate multi-class classification, SFID is applicable in this context, as illustrated in Figure 2, by employing multi-class classification with RandomForest.
For training the attribute classifier, we use the FairFace dataset, which includes seven racial categories: East Asian, Indian, Black, White, Middle Eastern, Latino Hispanic, and Southeast Asian. A detailed analysis and the corresponding results on multiple sensitive attributes are provided in Appendix F.
@@ 4.3 Dealing with Multiple Sensitive Attribute
Model Zero-shot Multi-class Classification
@@
Model Text-to-Image Retrieval
@@
||||Male-Female|||| Overall Composite
CLIP-CAP Baseline 34.57$\pm$0.83 25.41$\pm$0.73 2.20$\pm$1.81 2.10$\pm$0.70 3.24$\pm$1.61
DeAR 33.90$\pm$0.91 24.73$\pm$0.63 1.58$\pm$1.76 2.93$\pm$0.98 3.53$\pm$1.30
CLIP-clip 32.28$\pm$0.72 23.44$\pm$0.65 3.73$\pm$2.32 2.00$\pm$0.90 4.34$\pm$2.48
SFID (Ours) 32.08$\pm$0.78 23.74$\pm$0.69 2.16$\pm$2.03 2.07$\pm$1.03 3.12$\pm$1.82
BLIP Baseline 24.01$\pm$0.62 17.06$\pm$0.60 1.72$\pm$1.37 1.15$\pm$0.65 2.26$\pm$1.26
DeAR 21.76$\pm$0.59 15.51$\pm$0.47 2.62$\pm$1.84 1.07$\pm$0.63 2.84$\pm$2.13
CLIP-clip 23.74$\pm$0.54 16.96$\pm$0.54 2.29$\pm$1.67 1.15$\pm$0.65 2.59$\pm$1.81
SFID (Ours) 23.38$\pm$0.49 16.74$\pm$0.55 1.37$\pm$1.29 0.92$\pm$0.53 1.88$\pm$1.31
@@ 5.1 Implementation Detail
We use the FairFace dataset [22] for image inputs and the Bias-in-Bios dataset [12] for text inputs as our debiasing datasets, denoted as $X_{D}$. Each dataset is split into training and validation sets. A RandomForest classifier is applied to a 2D-shaped embedding representing all training samples, with low-confidence samples selected from the validation set. All hyperparameters and model settings for each baseline follow the default configurations provided in their respective open-source repositories. Detailed experimental settings, along with evaluation metrics and query datasets, are described in Section 3. We also adopt DeAR [34], CLIP-clip [38], and Prompt-Debias [11] as baseline comparison methods. A detailed discussion of these methods is provided in Appendix C.
For a fair comparison, we conduct 10 experiments using different subsets and report the mean and standard deviation for the text-to-image retrieval task. For zero-shot classification and image captioning, we employ 1000 bootstrapping iterations to calculate the confidence intervals. For text-to-image generation, images are generated using 10 random seeds. The mismatch rates are reported as the mean and standard deviation over 10 runs, while the Skew for the Neutral prompt is reported as a single value based on 10 runs. Further analysis of this evaluation metric can be found in Appendix E.
Model Mismatch Rate (Gender prompt) Neutral prompt
||||Male-Female|||| Overall Composite $Skew$
SDXL Baseline 3.87$\pm$2.23 2.35$\pm$1.22 4.42$\pm$2.57 83.25
DeAR 89.28$\pm$2.08 44.64$\pm$1.04 99.81$\pm$2.33 99.88
CLIP-clip 3.78$\pm$1.88 2.11$\pm$1.03 4.31$\pm$2.06 82.05
Prompt-Debias 39.72$\pm$6.83 42.53$\pm$3.85 58.49$\pm$3.64 82.77
SFID (LC) 1.69$\pm$0.72 0.96$\pm$0.42 1.97$\pm$0.67 81.57
SFID (HC) 1.54$\pm$1.14 0.84$\pm$0.71 1.74$\pm$1.57 81.57
CoDi Baseline 3.94$\pm$2.71 5.54$\pm$2.08 6.85$\pm$2.16 84.94
DeAR 5.63$\pm$2.84 5.42$\pm$1.10 8.05$\pm$3.00 86.14
CLIP-clip 4.73$\pm$2.22 5.00$\pm$1.39 7.01$\pm$1.53 84.58
Prompt-Debias 20.11$\pm$5.15 41.99$\pm$2.57 46.77$\pm$3.43 81.57
SFID (LC) 3.83$\pm$2.07 4.64$\pm$1.17 6.22$\pm$1.48 82.17
SFID (HC) 4.70$\pm$1.53 2.59$\pm$0.90 5.38$\pm$1.44 82.77
@@ 6 Conclusion
Acknowledgements
This work was partially supported by the EMBRIO Institute, contract #2120200, a National Science Foundation (NSF) Biology Integration Institute, Purdue’s Elmore ECE Emerging Frontiers Center, and NSF IIS #1955890, IIS #2146091, IIS #2345235.
@@ Appendix B Evaluation Metric for Image Captioning
METEOR measures the balance between precision and recall of n-grams in generated captions, incorporating synonyms. Let $P$ and $R$ be the precision and recall of matches between the generated caption and ground truth, including exact, synonym, and paraphrase matches: $\text{METEOR}=F_{\text{mean}}\cdot(1-\text{Pen})$ where $F_{\text{mean}}=\frac{10\cdot P\cdot R}{R+9\cdot P}$ is a harmonic mean, and $\text{Pen}=0.5\times\left(\frac{\text{number of chunks}}{\text{number of matches}}\right)^{3}$ is a penalty term where a chunk is a set of contiguous words in the generated captions that are in the reference. SPICE [3] focuses on the semantic content of captions by comparing sets of propositional semantic tuples extracted from candidate and reference captions. SPICE is the F1 score of precision and recall between the tuples of generated captions and ground truth.
Appendix C Comparison with Other Debiasing Approaches
As described in Section 1, two existing methods, DeAR [34] and CLIP-clip [38], utilize debiasing datasets with gender labels and pre-trained models to debias embeddings. While DeAR is specifically designed for image encoders and CLIP-clip for both image and text encoders, their methodologies could potentially be extended to text and image decoders, as well as other encoders. In contrast, Prompt-Debias [11] is designed primarily for text encoders, using pre-defined text prompts to achieve debiasing. These three approaches are selected for comparison in our analysis.
C.1 Debiasing with Additive Residuals (DeAR)
DeAR [34] involves a two-step training process. First, it trains an attribute classifier, $h_{c}$ using the frozen representation by minimizing the cross-entropy loss. Second, it trains an linear transformation $h_{a}$ for $Z_{D}$ to produce an additive feature to make a debiased representation $Z_{a}$, i.e., $Z_{a}=h_{a}(Z_{D})+Z_{D}$. The objective function for optimizing $h_{a}$
indicating the need to maintain the representation’s meaning while guiding the linear transformation $h_{a}$ to minimize the maximum probability of $h_{c}$ and maximize the cross-entropy loss of $h_{c}$. The goal is for the additive feature to make the representation ambiguous, placing it near the attribute classifier’s decision boundary. However, DeAR does not consider feature importance and applies adversarial training to all feature columns. Moreover, it requires several hyperparameters for the adversarial training, such as learning rate, weight decay, and the weights between loss functions ($\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$). This complexity may lead to inefficiencies in adversarial training, which is sensitive to hyperparameter settings, resulting in less effective translation of the feature, as shown in Figure 6. For DeAR training, all hyperparameters are set according to [34].
C.2 CLIP-clip
CLIP-clip [38] aims to identify features related to the sensitive attribute by measuring the mutual information between each column feature and the sensitive attribute. However, this approach assumes that each column independently contains meaningful information about the sensitive attribute, which can be overly restrictive. The dependency between feature maps in the neural networks’ representation output is well-documented [5, 36], supporting SFID’s capability in more effective feature selection.
Moreover, CLIP-clip is designed to clip the representation, reducing the dimension size of the embedding. The reduced embedding is not applicable as input for the decoder, which necessitates imputing the pruned features. Since CLIP-clip does not provide any suggestions for the imputed values, we adopted zero-value imputation across our experiments for CLIP-clip, where the number of pruned feature is set as $k=60$, showing best accuracy-fairness trade-off described in Appendix D.
C.3 Prompt-Debias
Prompt-Debias [11] seeks to mitigate bias in text embeddings by projecting out biased directions from the text input. However, Prompt-Debias is specifically designed for text-based tasks such as zero-shot classification and text-to-image generation. Unlike SFID, it lacks the flexibility to be applied across other components, limiting its effectiveness. This results in a performance constraint, as bias amplification can occur in the decoder or image encoder, components that Prompt-Debias cannot address. We include Prompt-Debias in our experiments for zero-shot classification, text-to-image retrieval, and text-to-image generation, excluding image captioning.
Appendix D The Impact of $k$ in CLIP-clip
@@
CLIP-clip [38] requires a hyperparameter $k$ to determine the number of pruned features. We selected $k$ for CLIP-clip based on empirical results with XVLM, as shown in Figures 7 and 8. The value $k=60$ performs best in text-to-image retrieval for the Flickr30K dataset, compromising slightly on accuracy and recall. Meanwhile, SFID demonstrates the best trade-off at $k=50$, corresponding to Appendix A.1.
@@
Appendix E Confidence Interval in Text-to-Image Generation
@@
On the other hand, a metric used in [11] measures gender discrepancy as follows:
@@
Appendix F Multiple Sensitive Attribute
@@
To mitigate this bias, SFID demonstrates its effectiveness even though the attributes in the training set and test set do not exactly match, i.e. race in FairFace and skin tone in FACET. We adopted an evaluation metric inspired by [16] and [14], which measures the maximum discrepancy across the sensitive attributes for each class. This metric is defined as follows:
@@
Appendix G Computational Resource